OpenStudy (anonymous):

Bottles of concentrated acids, such as phosphoric acid, H3PO4, often list the acid’s density and its mass percent. For example, concentrated solutions of phosphoric acid have a density of 1.69 g/mL are 85.5% H3PO4. What is the molarity of a solution of concentrated phosphoric acid?

OpenStudy (espex):

I believe this is how to do it, I am looking back in my notes to validate. Well, molarity is defined as moles/liter. \[H_3PO_4 = (3\times 1.008)+(1\times 30.974)+(4 \times 16.00)\rightarrow 97.998g\] Converting given to moles/L\[\frac{1 mole}{97.998g\space H_3PO_4} \times \frac{1000mL}{1Liter} \times \frac{1.69g}{1mL}\times .855=15.26\frac{moles}{Liter} (molarity)\]

OpenStudy (espex):

\[\frac{1mole\space H_3PO_4}{97.998g\space H_3PO_4} \times \frac {1000mL}{1 L} \times \frac{1.68g\space H_3PO_4 \times 0.855}{1mL}=14.74\frac{moles\space H_3PO_4}{Liter}\] The change is a result of taking 85.5% of the whole in the previous example.